OpenStudy (ebonyserrano):

A car travels at an average speed of 64 miles per hour. How long does it take to travel 368 miles?

OpenStudy (brooke..help00):

368 divided by 64?

jhonyy9 (jhonyy9):

use the rule of 3 simple 64 miles ----- 1 hour 368 miles ----- x hour -------------------- x = 368*1/64 = ?

jhonyy9 (jhonyy9):

hope helped and that is understandably sure right easy good luck bye bye

OpenStudy (ebonyserrano):

so 5 hours and 75 minutes? @jhonyy9

jhonyy9 (jhonyy9):

how you get this 75 minutes ? and do you know 1 hour how many minutes have ?

OpenStudy (brooke..help00):

That is correct.

OpenStudy (brooke..help00):

Isn't it @jhonyy9

jhonyy9 (jhonyy9):

brooke how can being correct ? please check it

OpenStudy (ebonyserrano):

would i change the 75?

jhonyy9 (jhonyy9):

than you not can making it directly so first of all you need convert the one hour equal how many minutes ?

jhonyy9 (jhonyy9):

64 miles ----- 1 hour 368 miles ----- x hour -------------------- x = 368*1/64 = ? and rewrite these above using just minutes in place of hour

OpenStudy (brooke..help00):

@jhonyy9 Msg me the correct answer so I can see where i messed up.

jhonyy9 (jhonyy9):

64 miles ------ 60 minutes 368 miles ----- x minutes -------------------- x = 368*60/64 = ? minutes fter this you need converting this result in hour and minutes

jhonyy9 (jhonyy9):

ebony do you understand it now ?

OpenStudy (ebonyserrano):

I multiplied 368 and 60 then divided by 64 and got 5.75

OpenStudy (brooke..help00):

Johnny message me pleas. I think we are all confused now.

jhonyy9 (jhonyy9):

ebony please check your result bc. not is right

OpenStudy (brooke..help00):

I got the same thing as her.

jhonyy9 (jhonyy9):

64 miles ----- 60 min. 368 miles ----- x min. -------------------- x = 368*60/64 = ?

jhonyy9 (jhonyy9):

x = 345 minutes

jhonyy9 (jhonyy9):

so after this you know that 1 hour equal 60 minutes x hour equal 345 min. ----------------------- x = 345*1/60 = 5 hours and 45 minutes

OpenStudy (brooke..help00):

@jhonyy9 Maybe you should double check your work. How did we get the same thing and you did not? What did you get?

jhonyy9 (jhonyy9):

this above wrote is the right answer

OpenStudy (ebonyserrano):

Thank you! I understand completely now!

jhonyy9 (jhonyy9):

ok. was my pleasure good luck bye bye

jhonyy9 (jhonyy9):

brooke do you understand it too ?

OpenStudy (brooke..help00):

Yeah I do too:)